OpenStudy (anonymous):

Solve 3^4x=72 Solve 2^x+7=125.93

OpenStudy (anonymous):

take a log

OpenStudy (anonymous):

thanks! then do I divide by 4?

OpenStudy (anonymous):

log(3^4x=72) 4x*log3 = log 72 4x* 0.477121255 =1.857332496 1.908485019*x=1.857332496 x=1.857332496/1.908485019 x=0.973197315

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

2^x+7=125.93 2^x=125.93-7 2^x=118.93 log(2^x=118.93) x*log2 = log 118.93 0.301029996 *x = 2.075291419 x=2.075291419/0.301029996 x=6.893968869